OpenStudy (anonymous):

cos(x+pi/3)-cos(x-pi/3)=-sqrt(3) the possible answers are: x= pi/2 x=2pi/3 +2pin x= pi/2 +2pin none of the above I attempted to use cos(u) - cos (v)= 2*sin((u+v/)2)*sin((u-v)/2) 2*sin(((x+pi/3)+(x-pi/3)/)2)*sin((x+pi/3)-(x-pi/3)/2) 2* sin(x) * sin( pi/3) 2* sqrt(3)/2* sin(x) sqrt(3)*sin(x) = -sqrt(3) sin(x)= -sqrt(3)/ sqrt(3) x= arcsin(-1) That would make x = -pi/2 correct?? Why is it not an option??

OpenStudy (anonymous):

Is it none of the above

OpenStudy (anonymous):

i ran -pi/2 though the equation and it isnt a solution, where did i mess up?

OpenStudy (anonymous):

Ah you need the formula \[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\]

OpenStudy (anonymous):

Keith... where would i use that formla

OpenStudy (anonymous):

From that you can break up \(\cos(x+\pi/3)\) and \(\cos(x-\pi/3)\)

OpenStudy (anonymous):

cos(x)cos(pi/3) - sin(x)sin(pi/3) - cos(x)cos(-pi/3) - sin(x)sin(-pi/3) like that?

OpenStudy (anonymous):

So you would have: \[\cos(x+\pi/3)-\cos(x-\pi/3) \\ =\cos(x)\cos(\pi/3) - \sin(x)\sin(\pi/3) - \cos(x)\cos(\pi/3) - \sin(x)\sin(\pi/3) \\ =-2\sin(x)\sin(\pi/3)=-\sqrt{3}\]

OpenStudy (anonymous):

And so with algebra you result in \[\sin(x)=\frac{\sqrt{3}}{2\sin(\pi/3)}\]

OpenStudy (anonymous):

shouldn't pi/3 be negative in the second half of the equation?

OpenStudy (anonymous):

No because the equation is \[cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\] So: \[cos(a-b)=cos(a+(-b))=\cos(a)\cos(-b)-\sin(a)\sin(-b)\] Since \(\cos(b)=\cos(-b)\) and \(\sin(-b)=-\sin(b)\) then: \[\cos(a-b)=\cos(a)\cos(b)-\sin(a)\times-\sin(b)=\cos(a)\cos(b)+\sin(a)\sin(b)\]

OpenStudy (anonymous):

I see, thanks :)

OpenStudy (anonymous):

Yeah so \(\pi/2\) is the correct answer